OpenStudy (nekoaisu):

Find the solution set. 13t^2+2=0

OpenStudy (campbell_st):

the solution will be a complex number start by subtracting 2 from both sides of the equation. then divide by 13 next take the square root to handle the square root of a negative use i^2 = -1 hope it helps

OpenStudy (nekoaisu):

Thank you!

OpenStudy (nekoaisu):

Is this answer correct? \[t = \frac{ \sqrt{2} }{ \sqrt{13} } i\]

OpenStudy (nekoaisu):

@campbell_st